Shift summary for Evening 11/7/99

Richard Ford (rickford@fnal.gov)
Mon, 8 Nov 1999 00:05:32 -0600 (CST)

Evening Shift, Nov. 7 1999

On Shift: E. Blucher, R. Ford

Hours of beam available at all intensities: 6.0 hours
Hours of beam available at nominal intensity: 6.0 hours
Hours of E799 data (special runs not included): 6.0 hours
Hours of special runs (laser scans, muon runs...): 0 hours

Accounting of lost time:

========================================================================
Runs ended this shift:

14922 BEAM_799 6.5 million events (data good after spill 4)

=======================================================================
Tapes taken to FCC this shift

none

========================================================================
Accesses during shift:

- adjust HA offsets following run 14922

========================================================================
Overview of shift: E799 datataking

Description of Shift:

14922 15:30 Inherit smooth running shift. 2 Hz still present.

16:56 146 Left bends trip. They are also going to reset QXR
to try and help the 2 Hz structure and minor beam
motion we've noticed

17:01 150 Beam returns

17:12 158 Beam disapperas MB7 interlock dropped and MuSep HV
problems. Experts being consulted.

18:00 At least an hour for experts to arrive.

Stop run 14922.

18:05 Rick F and Rick T make an access to adjust HA
LRS429 offsets.

149223 18:50 24 Beam returns. It appears that the MuSep problem
may not be as bad as originally thought.
==> beam structure is much better now.

The trigger rates do not seem to have been affected
by the changes in HA DC offsets.

121 Noticed the RC ADC distribution shifted slightly high
and also that the 4 trk/ 2trk ratio seems to be
drifting.

184 Quench at C4. Estimate 1 hour

23:40 230 Beam returns
========================================================================
Follow up Notes on Previous Shift